OpenStudy (anonymous):

Eigenvectors of this 3x3 matrix

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & 2 &-1\\ 2 &3 &-2 \\-1&-2&1\end{matrix}\right]\]

OpenStudy (turingtest):

did you get a characteristic polynomial yet?

OpenStudy (anonymous):

yep ive done it all but when i check iits wrong!! i have looked at this for over 30 mins!!! i write what i get \[\lambda = 5 \left(\begin{matrix}-1 \\ -2\\1\end{matrix}\right)\] \[\lambda = -1,-1, \left(\begin{matrix}-2 \\ -1\\0\end{matrix}\right),\left(\begin{matrix}0 \\ 1\\2\end{matrix}\right)\]

OpenStudy (anonymous):

sorry second matrix does not begin with -2 its 2 so 2 -1 0

OpenStudy (anonymous):

i have looked at my working out there is nothing wrong with it, but when i check with mathematica its wrong...

OpenStudy (anonymous):

no worries its correct, just check i wrote it wrong in matlab...